# Reciprocal theorem for calculating the flow rate of oscillatory channel flows

### v2

@@ Abstract
We demonstrate the use of the Lorentz reciprocal theorem in obtaining corrections to the steady flow rate due to flow oscillations in straight and nonuniform rigid channels. Starting from the unsteady Stokes equations, we derive the suitable reciprocity relation, assuming all quantities can be expressed as time-harmonic phasors. The auxiliary problem is the steady Hagen–Poiseuille flow solution, from which the reciprocal theorem allows us to calculate the first-order correction in the Womersley number to the steady flow rate. We also consider channels with nonuniform height in the flow-wise direction, in which case the flow-rate correction provides the leading-order effect of the interplay between the oscillations of the fluid flow and the given shape of the channel.
I Introduction
@@
Accurate modeling and design of such microsystems requires a quantitative understanding of the relationship between key integrated quantities, such as flow rate and pressure drop, in oscillatory flow. One approach to obtaining such relations is to solve the “complete” problem, meaning to solve the governing unsteady hydrodynamic equations. Another useful approach for finding integrated quantities (such as force, torque, or flow rate) is to exploit the Lorentz reciprocal theorem [8], which relates two different flow problems to each other (see, e.g., [9, 10] and the references therein). For example, in this way, a known solution of a “simpler” problem (termed the auxiliary problem) can be used to obtain the flow rate of a “harder” problem without the need for calculating its velocity and pressure field.
@@
This approach is particularly well-suited for low-Reynolds-number fluid mechanics applications such as those described above. For example, the reciprocal theorem has been used to predict the forces on and velocities of particles [11, 12, 13, 14, 15, 16, 17] and disks [11, 18] in oscillating flow fields. Although many applications of the reciprocal theorem are to three-dimensional flows, two-dimensional (2D) external oscillatory flows caused by the motion of immersed surfaces can also be handled [19, 20]. Furthermore, when combined with the lubrication approximation for slender flow geometries [21], the reciprocal theorem has also been used to obtain closed-form analytical expressions for flow rate–pressure drop relations of steady non-Newtonian flows through rigid [22, 23] and deformable [24, 25, 26] channels.
@@
However, despite all the progress made, to the best of our knowledge, no application of the reciprocal theorem to oscillatory channel flows has been presented to date. This observation motivates our study, which aims to show the reciprocal theorem’s versatility and simplicity for studying the coupling between oscillatory flow and channel shape.
@@
To this end, in Sec. II, we formulate the problem of low-Reynolds-number flow in two-dimensional channels, and construct a suitable reciprocal theorem relating an oscillatory flow to a steady flow. Next, in Sec. III, we introduce the lubrication approximation and show how the reciprocal theorem can be used to obtain the leading-order correction (in the Womersley number) to the flow rate due to flow oscillations. The calculation is then implemented for a straight 2D rigid channel in Sec. IV and for a nonuniform 2D rigid channel in Sec. V. Finally, Sec. VI concludes our study.
II Problem formulation and reciprocal theorem
@@
Consider the oscillatory flow of a Newtonian fluid in two 2D configurations, as shown in Fig. 1. The first configuration is a straight rigid channel (Fig. 1(a)). The second configuration is a nonuniform rigid channel with a spatially varying top wall (Fig. 1(b)). We assume that the axial length of the channels, $\ell$, is much greater than the channels’ height $h$, i.e., we consider slender geometries. The viscous fluid flow is driven by an imposed oscillatory pressure difference with amplitude $\Delta p$ and frequency $\omega$, resulting in a velocity field $\boldsymbol{v}$ and a pressure distribution $p$. Our goal is to determine the flow rate $q$, and specifically how it depends on $\omega$ (or, dimensionless groups involving $\omega$), without solving the oscillatory flow problem.
@@
$\boldsymbol{\nabla}\boldsymbol{\cdot}\boldsymbol{v}=0,\qquad\boldsymbol{\nabla}\boldsymbol{\cdot}\boldsymbol{\sigma}_{f}=\rho_{f}\frac{\partial\boldsymbol{v}}{\partial t}\qquad\text{with}\qquad\boldsymbol{\sigma}_{f}=-p\boldsymbol{I}+2\mu_{f}\boldsymbol{E},$ (1)
where $\boldsymbol{I}$ is the identity tensor, $\boldsymbol{E}=\tfrac{1}{2}[\boldsymbol{\nabla}\boldsymbol{v}+(\boldsymbol{\nabla}\boldsymbol{v})^{\top}]$ is the rate-of-strain tensor, $\mu_{f}$ is the fluid’s dynamic viscosity, and $\rho_{f}$ is the fluid’s density. To retain the unsteady term but neglect the convective inertia, as in Eq. (1), we introduce the Reynolds number ${\text{Re}}={\rho_{f}v_{c}h_{0}}/{\mu_{f}}$ and the Womersley number ${\text{Wo}}=h_{0}\sqrt{\rho_{f}\omega/\mu_{f}}$ [28, 27] and require that ${\text{Re}}\ll{\text{Wo}}^{2}$ (or, equivalently, $v_{c}/\omega\ll h_{0}$, i.e., the flow oscillation amplitude is much smaller than the channel height) [11]. Here, $v_{c}$ is the characteristic axial velocity scale and $h_{0}$ is either the rigid channel’s height or the inlet/outlet height of the nonuniform channel.
Figure 1: Schematic illustration of the two planar (2D) geometries considered: (a) a straight rigid channel and (b) a nonuniform rigid channel. The straight channel has a height $h_{0}$, which is also the inlet/outlet height of the nonuniform channel. The channels’ axial length is $\ell\gg h$. An oscillatory viscous fluid flow is driven by an imposed pressure difference, $\operatorname{Re}[\Delta p\,\mathrm{e}^{\mathrm{i}\omega t}]$, with the outlet $(z=\ell)$ pressure setting the gage. We are interested in determining the resulting flow rate $q$.
@@
$\boldsymbol{v}(y,z,t)=\operatorname{Re}[\boldsymbol{v}_{a}(y,z)\mathrm{e}^{\mathrm{i}\omega t}],\qquad p(y,z,t)=\operatorname{Re}[p_{a}(y,z)\mathrm{e}^{\mathrm{i}\omega t}],\qquad\boldsymbol{\sigma}_{f}(y,z,t)=\operatorname{Re}[\boldsymbol{\sigma}_{f,a}(y,z)\mathrm{e}^{\mathrm{i}\omega t}],$ (2)
where the quantities with an ‘$a$’ subscript are complex-valued amplitudes. Using Eq. (2), the linearity of the continuity and momentum equations (1), and leaving the $\operatorname{Re}[\,\cdot\,]$ understood, we obtain
$\boldsymbol{\nabla}\boldsymbol{\cdot}\boldsymbol{v}_{a}=0,\qquad\boldsymbol{\nabla}\boldsymbol{\cdot}\boldsymbol{\sigma}_{f,a}=\rho_{f}\omega\mathrm{i}\boldsymbol{v}_{a}\qquad\text{with}\qquad\boldsymbol{\sigma}_{f,a}=-p_{a}\boldsymbol{I}+2\mu_{f}\boldsymbol{E}_{a}.$ (3)
Now, let $\hat{\boldsymbol{v}}$, $\hat{p}$, and $\hat{\boldsymbol{\sigma}}_{f}$ denote, respectively, the velocity, pressure, and stress fields corresponding to the steady solution in the same channel, satisfying
$\boldsymbol{\nabla}\boldsymbol{\cdot}\hat{\boldsymbol{v}}=0,\qquad\boldsymbol{\nabla}\boldsymbol{\cdot}\hat{\boldsymbol{\sigma}}_{f}=\boldsymbol{0}\qquad\text{with}\qquad\hat{\boldsymbol{\sigma}}_{f}=-\hat{p}\boldsymbol{I}+2\mu_{f}\hat{\boldsymbol{E}}$. (4)
Forming the product of the momentum balances in Eqs. (3) and (4), with $\hat{\boldsymbol{v}}$ and $\boldsymbol{v}_{a}$, respectively, and using the usual vector identities, incompressibility, and the symmetry of the stress tensor [10, 22], we obtain
$\boldsymbol{\nabla}\boldsymbol{\cdot}(\boldsymbol{\sigma}_{f,a}\boldsymbol{\cdot}\hat{\boldsymbol{v}})-\boldsymbol{\sigma}_{f,a}\boldsymbol{:}\hat{\boldsymbol{E}}$ $=\rho_{f}\omega\mathrm{i}\boldsymbol{v}_{a}\boldsymbol{\cdot}\hat{\boldsymbol{v}}$, (5a)
$\boldsymbol{\nabla}\boldsymbol{\cdot}(\hat{\boldsymbol{\sigma}}_{f}\boldsymbol{\cdot}\boldsymbol{v}_{a})-\hat{\boldsymbol{\sigma}}_{f}\boldsymbol{:}\boldsymbol{E}_{a}$ $=0.$ (5b)
@@
$\boldsymbol{\nabla}\boldsymbol{\cdot}(\boldsymbol{\sigma}_{f,a}\boldsymbol{\cdot}\hat{\boldsymbol{v}})-\boldsymbol{\nabla}\boldsymbol{\cdot}(\hat{\boldsymbol{\sigma}}_{f}\boldsymbol{\cdot}\boldsymbol{v}_{a})=\rho_{f}\omega\mathrm{i}\boldsymbol{v}_{a}\boldsymbol{\cdot}\hat{\boldsymbol{v}}$. (6)
@@
$\int_{S_{0}}\boldsymbol{n}\boldsymbol{\cdot}\boldsymbol{\sigma}_{f,a}\boldsymbol{\cdot}\hat{\boldsymbol{v}}\,\mathrm{d}S-\int_{S_{0}}\boldsymbol{n}\boldsymbol{\cdot}\hat{\boldsymbol{\sigma}}_{f}\boldsymbol{\cdot}\boldsymbol{v}_{a}\,\mathrm{d}S+\int_{S_{\ell}}\boldsymbol{n}\boldsymbol{\cdot}\boldsymbol{\sigma}_{f,a}\boldsymbol{\cdot}\hat{\boldsymbol{v}}\,\mathrm{d}S-\int_{S_{\ell}}\boldsymbol{n}\boldsymbol{\cdot}\hat{\boldsymbol{\sigma}}_{f}\boldsymbol{\cdot}\boldsymbol{v}_{a}\,\mathrm{d}S=\rho_{f}\omega\mathrm{i}\int_{\mathcal{V}}\boldsymbol{v}_{a}\boldsymbol{\cdot}\hat{\boldsymbol{v}}\,\mathrm{d}\mathcal{V},$ (7)
similar to [17]. Here, $\boldsymbol{n}$ is the unit outward normal to $S_{0,\ell}$. Note that the integrals over the bottom and top walls of the channel vanish because of the no-slip and no-penetration boundary conditions, $\hat{\boldsymbol{v}}=\boldsymbol{v}_{a}=\boldsymbol{0}$, there.
III Nondimensionalization and lubrication approximation
@@
$T=\frac{t}{t_{c}},\quad Y=\frac{y}{h_{0}},\quad Z=\frac{z}{\ell},\quad V_{y}=\frac{v_{y}}{\epsilon v_{c}},\quad V_{z}=\frac{v_{z}}{v_{c}},\quad Q=\frac{q}{q_{c}},\quad P=\frac{p}{p_{c}},\quad H=\frac{h}{h_{0}},$ (8)
where $p_{c}=\mu_{f}v_{c}\ell/h_{0}^{2}$ is the characteristic lubrication pressure scale, $t_{c}=1/\omega$ is the characteristic time scale, and $\epsilon=h_{0}/\ell$ is the slenderness ratio of the channel, assumed to be small, $\epsilon\ll 1$, which is lubrication approximation [27].
We consider only planar geometries and analyze the problems per unit width. The flow is driven by an imposed pressure drop, which leads to $p_{c}=\Delta p\stackrel{{\scriptstyle\text{def}}}{{=}}p|_{z=0}-p|_{z=\ell}$ so that the characteristic axial velocity scale is $v_{c}=h_{0}^{2}\Delta p/(\mu_{f}\ell)$ and, consequently, the characteristic flow rate scale is $q_{c}=h_{0}v_{c}=h_{0}^{3}\Delta p/(\mu_{f}\ell)$. The regime of imposed flow rate is considered in Appendix A.
@@
$\left.\boldsymbol{n}\boldsymbol{\cdot}\boldsymbol{\sigma}_{f,a}\boldsymbol{\cdot}\hat{\boldsymbol{v}}\right|_{z=0,\,\ell}$ $\simeq\mp\frac{\mu_{f}v_{c}^{2}\ell}{h_{0}^{2}}\left[-P_{a}\hat{V}_{z}\right]_{Z=0,\,1},$ (9a)
$\left.\boldsymbol{n}\boldsymbol{\cdot}\hat{\boldsymbol{\sigma}}_{f}\boldsymbol{\cdot}\boldsymbol{v}_{a}\right|_{z=0,\,\ell}$ $\simeq\mp\frac{\mu_{f}v_{c}^{2}\ell}{h_{0}^{2}}\left[-\hat{P}V_{z,a}\right]_{Z=0,\,1},$ (9b)
@@
$\int_{0}^{H(0)}\left[P_{a}\hat{V}_{z}\right]_{Z=0}\,\mathrm{d}Y-\int_{0}^{H(0)}\left[\hat{P}V_{z,a}\right]_{Z=0}\,\mathrm{d}Y=\mathrm{i}{\text{Wo}}^{2}\int_{0}^{1}\int_{0}^{H(Z)}V_{z,a}\hat{V}_{z}\,\mathrm{d}Y\mathrm{d}Z.$ (10)
Noting that $P_{a}=P_{a}(Z)$ and $\hat{P}=\hat{P}(Z)$ under the lubrication approximation [27], and defining $\int_{0}^{H(Z)}\hat{V}_{z}\mathrm{d}Y\stackrel{{\scriptstyle\text{def}}}{{=}}\hat{Q}$ and $\int_{0}^{H(Z)}V_{z,a}\mathrm{d}Y\stackrel{{\scriptstyle\text{def}}}{{=}}Q_{a}$, which are necessarily constant, Eq. (10) yields the reciprocal theorem,
$\Delta P_{a}\hat{Q}=\Delta\hat{P}Q_{a}+\mathrm{i}{\text{Wo}}^{2}\int_{0}^{1}\int_{0}^{H(Z)}V_{z,a}\hat{V}_{z}\,\mathrm{d}Y\mathrm{d}Z,$ (11)
relating the oscillatory flow (‘$a$’ subscripts) of a Newtonian fluid in a 2D channel to its steady counterpart (‘hats’). Note that Eq. (11), which is the key result of this work, is valid for both configurations in Fig. 1.
Equation (11) clearly shows that the dimensionless amplitude of the flow rate (or pressure drop) depends on $V_{z,a}$, and thus, generally, requires solving the complete oscillatory flow problem. However, in the weakly oscillatory limit, corresponding to ${\text{Wo}}^{2}\ll 1$, we can apply the reciprocal theorem (11) to obtain the leading-order oscillatory correction to the flow rate using only the steady solution of the problem, as also observed in [15, 17]. To this end, we expand the amplitude of velocity into a perturbation series in the dimensionless parameter ${\text{Wo}}^{2}\ll 1$,
$V_{z,a}=\hat{V}_{z}+{\text{Wo}}^{2}V_{z,a,1}+\mathrm{O}({\text{Wo}}^{4}),$ (12)
where $V_{z,a,1}$ is the first-order correction to the phasor amplitude of the axial velocity due to the oscillatory flow, and $V_{z,a,0}\equiv\hat{V}_{z}$ because the steady flow is the zero-Womersley-number limit of the oscillatory one (recall Eqs. (1) and (3)). The remaining calculations are thus all accurate to $\mathrm{O}({\text{Wo}}^{2})$.
IV Oscillatory flow in a straight rigid channel
Consider the straight rigid channel, shown in Fig. 1(a), such that $H=1$. The dimensionless solution of the steady flow in this configuration, in a pressure-drop-controlled regime, is
$\hat{Q}=\frac{\Delta\hat{P}}{12},\qquad\hat{V}_{z}(Y)=\frac{1}{2}\Delta\hat{P}(1-Y)Y.$ (13)

### v3

@@ Abstract
We demonstrate the use of the Lorentz reciprocal theorem in obtaining corrections to the steady flow rate due to flow oscillations in rigid channels. Starting from the unsteady Stokes equations, we derive the suitable reciprocity relation, assuming all quantities can be expressed as time-harmonic phasors. The auxiliary problem is the steady Hagen–Poiseuille flow solution, from which the reciprocal theorem allows us to calculate the first-order correction in the Womersley number to the steady flow rate in a straight rigid channel. We also consider nonuniform channels, specifically with variable height in the flow-wise direction, in which case the flow rate correction provides the leading-order effect of the interplay between the oscillations of the fluid flow and the given shape of the channel.
keywords:
Oscillatory channel flow , Reciprocal theorem , Flow rate
[…]
[PUME]organization=School of Mechanical Engineering, Purdue University,city=West Lafayette, postcode=47907, state=Indiana, country=USA
[…]
[Technion]organization=Faculty of Mechanical Engineering, Technion – Israel Institute of Technology,city=Haifa, postcode=3200003, country=Israel
1 Introduction
@@
Accurate modeling and design of such microsystems requires a quantitative understanding of the relationship between key integrated quantities, such as flow rate and pressure drop, in oscillatory flow. One approach to obtaining such relations is to solve the “complete” problem, meaning to solve the governing unsteady hydrodynamic equations. Another useful approach for finding integrated quantities (such as force, torque, or flow rate) is to exploit the Lorentz reciprocal theorem [8], which relates two different flow problems to each other (see, e.g., [9, 10] and the references therein). For example, in this way, a known solution of a “simpler” problem (termed the auxiliary problem) can be used to obtain the flow rate of a “harder” problem without the need for calculating its velocity and pressure fields.
@@
This approach is particularly well-suited for low-Reynolds-number fluid mechanics applications such as those described above. For example, the reciprocal theorem has been used to predict the forces on and velocities of particles [11, 12, 13, 14, 15, 16, 17] and disks [11, 18] in oscillating flow fields. Although many applications of the reciprocal theorem are to three-dimensional flows, two-dimensional (2D) external oscillatory flows caused by the motion of immersed surfaces can also be handled [19, 20]. In the present context, 2D flows are an idealization of 3D flows that are unconfined (or of large extent) in the spanwise direction perpendicular to the flow. Then, the lubrication approximation for slender flow geometries [21] suggests that the velocity varies transversely and the pressure varies axially, and the reciprocal theorem has been used to obtain closed-form analytical expressions for flow rate–pressure drop relations of steady non-Newtonian flows through rigid [22, 23] and deformable [24, 25, 26] channels.
@@
However, despite all the progress made, to the best of our knowledge, no application of the reciprocal theorem to oscillatory channel flows has been presented to date. This observation motivates our study, which aims to show the reciprocal theorem’s versatility and simplicity for studying the coupling between flow oscillations and channel shape.
@@
To this end, in Sec. 2, we formulate the problem of low-Reynolds-number flow in two-dimensional channels, and construct a suitable reciprocal theorem relating an oscillatory flow to a steady flow. Next, in Sec. 3, we introduce the lubrication approximation and show how the reciprocal theorem can be used to obtain the leading-order correction (in the Womersley number) to the flow rate due to flow oscillations. The calculation is then implemented for a straight 2D rigid channel in Sec. 4 and for a nonuniform 2D rigid channel, specifically with variable height in the flow-wise direction, in Sec. 5. Finally, Sec. 6 concludes our study.
2 Problem formulation and reciprocal theorem
@@
Consider the incompressible oscillatory flow of a Newtonian fluid in the two 2D configurations shown in Fig. 1. The first configuration is a straight rigid channel of uniform height (Fig. 1(a)). The second configuration is a nonuniform rigid channel with variable height in the flow-wise direction (Fig. 1(b)). We assume that the axial length of the channels, $\ell$, is much greater than the channels’ height $h$, i.e., we consider slender geometries. The viscous fluid flow is driven by an imposed oscillatory pressure difference with amplitude $\Delta p$ and frequency $\omega$, resulting in a velocity field $\bm{v}$ and a pressure distribution $p$. Our goal is to determine the flow rate $q$, and specifically how it depends on $\omega$ (or, dimensionless groups involving $\omega$), without solving the oscillatory flow problem.
@@
$\bm{\nabla}\bm{\cdot}\bm{v}=0,\quad\bm{\nabla}\bm{\cdot}\bm{\sigma}=\rho\frac{\partial\bm{v}}{\partial t},\quad\bm{\sigma}=-p\bm{I}+2\mu\bm{E},$ (1)
where $\bm{I}$ is the identity tensor, $\bm{E}=\tfrac{1}{2}[\bm{\nabla}\bm{v}+(\bm{\nabla}\bm{v})^{\top}]$ is the rate-of-strain tensor, $\mu$ is the fluid’s dynamic viscosity, and $\rho$ is the fluid’s density. To retain the unsteady term but neglect the convective inertia, as in Eq. (1), we introduce the Reynolds number $\mathrm{Re}={\rho v_{c}h_{0}}/{\mu}$ and the Womersley number $\mathrm{Wo}=h_{0}\sqrt{\rho\omega/\mu}$ [28, 27] and require that $\mathrm{Re}\ll\mathrm{Wo}^{2}$ (or, equivalently, $v_{c}/\omega\ll h_{0}$, i.e., the flow oscillation amplitude is much smaller than the channel height) [11]. Here, $v_{c}$ is the characteristic axial velocity scale, and $h_{0}$ is either the rigid channel’s height or the inlet/outlet height of the nonuniform channel.
Figure 1: Schematic illustration of the two planar (2D) geometries considered: (a) a straight rigid channel and (b) a nonuniform rigid channel, specifically with variable height in the flow-wise direction. The straight channel has a height $h_{0}$, which is also the inlet/outlet height of the nonuniform channel. The integration domain $\mathcal{V}$ for the reciprocal theorem is the shaded interior of the channels. Its boundary is marked with bright dotted curves; $S_{0}$ and $S_{\ell}$ denote the inlet and outlet surfaces, respectively. The channels’ axial length is $\ell\gg h$. An oscillatory viscous fluid flow is driven by an imposed pressure difference, $\Re[\Delta p\,\mathrm{e}^{\mathrm{i}\omega t}]$, with the outlet $(z=\ell)$ pressure setting the gage. The fluid velocity satisfies the no-slip condition along the stationary top and bottom walls of the channels (i.e., $\bm{v}=\bm{0}$ there). We are interested in determining the resulting flow rate $q$.
@@
$\bm{v}(y,z,t)$ $=\Re[\bm{v}_{a}(y,z)\mathrm{e}^{\mathrm{i}\omega t}],$ (2a)
$p(y,z,t)$ $=\Re[p_{a}(y,z)\mathrm{e}^{\mathrm{i}\omega t}],$ (2b)
$\bm{\sigma}(y,z,t)$ $=\Re[\bm{\sigma}_{a}(y,z)\mathrm{e}^{\mathrm{i}\omega t}],$ (2c)
where the quantities with an ‘$a$’ subscript are complex-valued amplitudes. Using Eqs. (2), the linearity of the continuity and momentum equations (1), and leaving the $\Re[\,\cdot\,]$ understood, we obtain
$\bm{\nabla}\bm{\cdot}\bm{v}_{a}=0,\quad\bm{\nabla}\bm{\cdot}\bm{\sigma}_{a}=\rho\omega\mathrm{i}\bm{v}_{a},\quad\bm{\sigma}_{a}=-p_{a}\bm{I}+2\mu\bm{E}_{a}.$ (3)
Now, let $\hat{\bm{v}}$, $\hat{p}$, and $\hat{\bm{\sigma}}$ denote, respectively, the velocity, pressure, and stress fields corresponding to the steady solution in the same channel, satisfying
$\bm{\nabla}\bm{\cdot}\hat{\bm{v}}=0,\quad\bm{\nabla}\bm{\cdot}\hat{\bm{\sigma}}=\bm{0},\quad\hat{\bm{\sigma}}=-\hat{p}\bm{I}+2\mu\hat{\bm{E}}$. (4)
Forming the product of the momentum balances in Eqs. (3) and (4), with $\hat{\bm{v}}$ and $\bm{v}_{a}$, respectively, and using the usual vector identities, incompressibility, and the symmetry of the stress tensor [10, 22], we obtain
$\bm{\nabla}\bm{\cdot}(\bm{\sigma}_{a}\bm{\cdot}\hat{\bm{v}})-\bm{\sigma}_{a}\bm{:}\hat{\bm{E}}$ $=\rho\omega\mathrm{i}\bm{v}_{a}\bm{\cdot}\hat{\bm{v}}$, (5a)
$\bm{\nabla}\bm{\cdot}(\hat{\bm{\sigma}}\bm{\cdot}\bm{v}_{a})-\hat{\bm{\sigma}}\bm{:}\bm{E}_{a}$ $=0.$ (5b)
@@
$\bm{\nabla}\bm{\cdot}(\bm{\sigma}_{a}\bm{\cdot}\hat{\bm{v}})-\bm{\nabla}\bm{\cdot}(\hat{\bm{\sigma}}\bm{\cdot}\bm{v}_{a})=\rho\omega\mathrm{i}\bm{v}_{a}\bm{\cdot}\hat{\bm{v}}$. (6)
@@
$\int_{S_{0}}\bm{n}\bm{\cdot}\bm{\sigma}_{a}\bm{\cdot}\hat{\bm{v}}\,\mathrm{d}S-\int_{S_{0}}\bm{n}\bm{\cdot}\hat{\bm{\sigma}}\bm{\cdot}\bm{v}_{a}\,\mathrm{d}S\\ +\int_{S_{\ell}}\bm{n}\bm{\cdot}\bm{\sigma}_{a}\bm{\cdot}\hat{\bm{v}}\,\mathrm{d}S-\int_{S_{\ell}}\bm{n}\bm{\cdot}\hat{\bm{\sigma}}\bm{\cdot}\bm{v}_{a}\,\mathrm{d}S=\rho\omega\mathrm{i}\int_{\mathcal{V}}\bm{v}_{a}\bm{\cdot}\hat{\bm{v}}\,\mathrm{d}\mathcal{V}.$ (7)
Here, $\bm{n}$ is the unit outward normal to $S_{0,\ell}$. Note that the integrals over the bottom and top walls of the channel vanish because of the no-slip boundary conditions, $\hat{\bm{v}}=\bm{v}_{a}=\bm{0}$, there. In passing, we note that Eq. (7) is similar to Eq. (3.13) of Zhang et al. [17], who used the reciprocal theorem to obtain the force on an immersed sphere oscillating near a wall.
3 Nondimensionalization and lubrication approximation
@@
$T=\frac{t}{t_{c}},\quad Y=\frac{y}{h_{0}},\quad Z=\frac{z}{\ell},\quad V_{y}=\frac{v_{y}}{\epsilon v_{c}},\quad V_{z}=\frac{v_{z}}{v_{c}},\\ Q=\frac{q}{q_{c}},\quad P=\frac{p}{p_{c}},\quad H=\frac{h}{h_{0}},$ (8)
where $p_{c}=\mu v_{c}\ell/h_{0}^{2}$ is the characteristic lubrication pressure scale, $t_{c}=1/\omega$ is the characteristic time scale, and $\epsilon=h_{0}/\ell$ is the slenderness ratio of the channel, assumed to be small, i.e., $\epsilon\ll 1$; this assumption is the lubrication approximation [27].
We consider only planar geometries and analyze the problems per unit width. The flow is driven by an imposed pressure drop, which leads to $p_{c}=\Delta p\stackrel{{\scriptstyle\text{def}}}{{=}}p|_{z=0}-p|_{z=\ell}$ so that the characteristic axial velocity scale is $v_{c}=h_{0}^{2}\Delta p/(\mu\ell)$ and, consequently, the characteristic flow rate scale is $q_{c}=h_{0}v_{c}=h_{0}^{3}\Delta p/(\mu\ell)$. The regime of imposed flow rate is considered in A.
@@
$\left.\bm{n}\bm{\cdot}\bm{\sigma}_{a}\bm{\cdot}\hat{\bm{v}}\right|_{z=0,\,\ell}$ $\simeq\mp\frac{\mu v_{c}^{2}\ell}{h_{0}^{2}}\left[-P_{a}\hat{V}_{z}\right]_{Z=0,\,1},$ (9a)
$\left.\bm{n}\bm{\cdot}\hat{\bm{\sigma}}\bm{\cdot}\bm{v}_{a}\right|_{z=0,\,\ell}$ $\simeq\mp\frac{\mu v_{c}^{2}\ell}{h_{0}^{2}}\left[-\hat{P}V_{z,a}\right]_{Z=0,\,1},$ (9b)
@@
$\int_{0}^{H(0)}\left[P_{a}\hat{V}_{z}\right]_{Z=0}\,\mathrm{d}Y-\int_{0}^{H(0)}\left[\hat{P}V_{z,a}\right]_{Z=0}\,\mathrm{d}Y\\ =\mathrm{i}\mathrm{Wo}^{2}\int_{0}^{1}\int_{0}^{H(Z)}V_{z,a}\hat{V}_{z}\,\mathrm{d}Y\mathrm{d}Z.$ (10)
Noting that $P_{a}=P_{a}(Z)$ and $\hat{P}=\hat{P}(Z)$ under the lubrication approximation [27], and defining the flow rates, $\int_{0}^{H(Z)}\hat{V}_{z}\,\mathrm{d}Y\stackrel{{\scriptstyle\text{def}}}{{=}}\hat{Q}$ and $\int_{0}^{H(Z)}V_{z,a}\,\mathrm{d}Y\stackrel{{\scriptstyle\text{def}}}{{=}}Q_{a}$, which are necessarily constant, Eq. (10) yields the reciprocal theorem,
$\Delta P_{a}\hat{Q}=\Delta\hat{P}Q_{a}+\mathrm{i}\mathrm{Wo}^{2}\int_{0}^{1}\int_{0}^{H(Z)}V_{z,a}\hat{V}_{z}\,\mathrm{d}Y\mathrm{d}Z,$ (11)
relating the oscillatory flow (‘$a$’ subscripts) of a Newtonian fluid in a 2D channel to its steady counterpart (‘hats’). Note that Eq. (11), which is the key result of this work, is valid for both oscillatory flow in a straight rigid channel and oscillatory flow in a nonuniform rigid channel with variable height in the flow-wise direction—the two configurations shown in Fig. 1.
Equation (11) clearly shows that the dimensionless amplitude of the flow rate (or pressure drop) depends on $V_{z,a}$, and thus, generally, requires solving the complete oscillatory flow problem. However, in the weakly oscillatory limit, corresponding to $\mathrm{Wo}^{2}\ll 1$, we can apply the reciprocal theorem (11) to obtain the leading-order oscillatory correction to the flow rate using only the steady solution of the problem, as also observed in [15, 17]. To this end, we expand the amplitude of the velocity into a perturbation series in the dimensionless parameter $\mathrm{Wo}^{2}\ll 1$,
$V_{z,a}=\hat{V}_{z}+\mathrm{Wo}^{2}V_{z,a,1}+\mathrm{O}(\mathrm{Wo}^{4}),$ (12)
where $V_{z,a,1}$ is the first-order correction to the phasor amplitude of the axial velocity due to the oscillatory flow, and $V_{z,a,0}\equiv\hat{V}_{z}$ because the steady flow is the zero-Womersley-number limit of the oscillatory one (recall Eqs. (1) and (3)). The remaining calculations are thus all accurate to $\mathrm{O}(\mathrm{Wo}^{2})$.
4 Oscillatory flow in a straight rigid channel
Consider the straight rigid channel, shown in Fig. 1(a), for which $H=1$. The dimensionless solution of the steady flow in this configuration, in a pressure-drop-controlled regime, is
$\hat{Q}=\frac{\Delta\hat{P}}{12},\quad\hat{V}_{z}(Y)=\frac{1}{2}\Delta\hat{P}(1-Y)Y.$ (13)